Solve and Check.
17x^2-10x=0

Answer by jim_thompson5910(35256)

17x^2-10x=0

x(17x-10)=0

x = 0 or 17x-10 = 0

x = 0 or 17x = 10

x = 0 or x = 10/17

So the solutions are x = 0 or x = 10/17

I'll leave the check for you to do.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------